Widget Co. was founded at the start of 2016 by issuing 10M shares at $187.50 per share and $1,200M of long term debt. All data in the table, except actual price per share, is in millions. No assets were sold during the first two years of operation. Remember when and where to include marketable securities and the current portion of long term debt in your intermediate calculations. Use this information and the table below to answer problems 20-3O, 2016 250 125 100 100 300 300 3000 200 2800 400 2017 300 150 Net Income Dividends Cash Marketable Sec. Acct. Rec. Inventory Gross PPE Accumulated Dep. Net PPE Acct. Payable CP LT Debt LT Debt 120 300 360 3600 500 3100 400 100 1400 Shares Out P/sh 1200 100 20.00 22.00 28. Between December 2016 and December 2017, Widget Companys market capitalization: a. Increased by $332M b. Increased by $200M c. Increased by $156M d. Decreased by $1 56M e. Decreased by $200M 29. At the end of 2016, Widget Companys P/E is closest to: α, 72 b. 8.0 c. 14.4 d. 15.0 e. 16.0 30. During 2017, Widget Company HINT: BOY SE +NI - DIVS +Equity Issued- Equity RepurchasedEOY SE 23 a. Net borrowing was $200M and shares outstanding were repurchased at a price of $24 per share b. Net borrowing was $200M and the company issued shares at a price of $25 per share c. Net borrowing was $300M and shares outstanding were repurchased at a price of $24 per share d. Net borrowing was $300M and shares outstanding were repurchased at a price of $25 per share e. Net borrowing was $300M and the company issued shares at a price of $24 per share
Answer #1

28)

Time Outstanding shares Price/share Market cap.
Dec 2006 100 20 2000
Dec 2007 98 22 2156

2156 - 2000 = increase of 156 million - option c.

29)

price/Share 20
Earnings / share 250/100 = 2.5
P/E 20/2.5 = 8

option b.